sinx +cosx=k find value of sin2x

sinx +cosx=k
find value of sin2x

Grade:7

1 Answers

Shubham Chakraverty
10 years ago
(sin x + cos x)^2=k^2 , 1 + sin 2x = k^2 , sin 2x = k^2 -1